OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

OpenStudy (anonymous):

is it mx+b?

OpenStudy (anonymous):

slope?

OpenStudy (anonymous):

is that a choice?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what r?

OpenStudy (anonymous):

what are you choices for your problem

OpenStudy (anonymous):

no choices

OpenStudy (anonymous):

slope = MX+B are you in slope?

OpenStudy (anonymous):

it should give you choices

OpenStudy (anonymous):

The solution is \[x=6\] and \[y=12\] To find this, multiply the second equation by \[-1\] Then add the two equations. This gets rid of the \[y\] Solve for \[x\] and then sub this value into one of the equations to determine \[y\]

OpenStudy (anonymous):

like point slope which i know it isn't, slope intercept form, slope

OpenStudy (anonymous):

Alternatively, instead of using elimination, the above method, You can use Substitution. Given y = x +6. Substitute this for y in the next equation. (x+6) = 2x. Solve for x. 6 = x now plug back in to find y. y = x+6. so y = 6+6. y = 12. Therefore we have (6,12) as the solution

OpenStudy (anonymous):

they have it so eacortez pick the one of the guys or girls with the best response if there awnsewr helps. see ya bro